Saturday, 6 May 2017

Maintenance Update (clean, spring, oil, weights)

Wednesday 3rd
At last, the engine stood still long enough for us to address some of the issues!

Brian seemed to be in “management” role, as he supervised Tom, Alex and Gwen who were cleaning various bits on the loco.

Brian also gave a brief report on the recent visit to Llangollen, which is one of the places under consideration for the boiler overhaul in 2020. It was interesting to note that they are already booked up for boiler work until the beginning of 2020! Brian also examined the leaking ash pan sprinkler, where it became apparent that the delivery pipe had sheared where it passes into the ash pan. This may have been the result of expansion/contraction stress.

Gil told me that Tom and Roy from the Loco Dept had cleaned out the drain cocks (which had allegedly stuck at one point last week).

Bruce examined the spring that has been reported as possibly broken. Clive had also taken a look, and it does not appear to be broken, but we will keep an eye on it. Bruce also investigated the alleged oil leak from the pony truck. His view is that it was most likely over-filled, and it is designed to dribble oil down the horn guides, which may have triggered the report. The oil pot was full, and no sign of leakage.

Bruce had another bash (literally) at the balance weights on the driving wheel to make sure they are as tight as is possible.

Also, he’d been calculating the adjustment required on the safety valves to get blow-off closer to the 225 mark (as they currently lift at 210).

Clive replaced the broken spring in the flange lubricator.

John G was painting rail chairs all day: one blue (GER), one red (MR) and eight green.

I was also working on the boot scraper production line for much of the day, apart from assisting PWay with the cutting of some rail for a customer.

Here is the list of issues:
1 J cocks are stuck [A last-year’s issue: We never did find out which J cocks were stuck …
I decided to cross it off unless/until a clearer report is given in the future]
2 LH front & rear top mud hole doors & right-hand top blowing [Marked as “Monitor”]
7 Safety valves lift at 215 lbs [We will verify and adjust at next wash-out]
8 Top front left hand mud hole door weeping [Repeat of 2]
9 Top front right hand mud hole door weeping [Repeat of 2]
12 RH piston gland packing blowing [Gil reckons that this is not a significant leak]
13 LH pony axlebox losing oil but cool at the end of the day [Suspect driver mistaken]
14 Mud hole door RHS of firebox leaking [Repeat of 2]
16 Suspect broken leaf in LH trailing wheel [spring], 3rd or 4th leaf down. [Not broken, but will
17 RH trailing flange lubricator spring broken. [Repaired]

Saturday 6th
2807 was on the Fire & Drive on Friday, and continued in service today. Only John T and I were at
Todders, and there was little to do apart from boot scraper production.

Gilbert popped in, but toddled off to Winchcombe to help Fred on the siphon restoration.

Sunday is 2807’s last service until the gala (27th - 29th May). There are no new issues logged.


No comments: